OpenStudy (anonymous):

1 MEDAL HELP PLEASE AS SOON AS POSSIBLE. Select the best possible first step to solving the system by first eliminating the x variable. 3x - 9y = 6 2x - 11y = 6

Directrix (directrix):

@maaryliin Please post the steps >>>Select the best possible first step

OpenStudy (anonymous):

multiplying equation 1 by -2/3 or multipling eq 2 by -3/2

Directrix (directrix):

1) 3x - 9y = 6 Multiply by -2/3 -2/3 * 3x -2/3 * (-9)y = (-2/3)*6 1* -2x + 6y = -4 Compare 1* with Equation 2 1*) -2x + 6y = -4 2) 2x - 11y = 6 -------------------- Notice that the x terms -2x and 2x add to 0x and the x-term is eliminated.

Directrix (directrix):

But, in applying the other option, multipling eq 2 by -3/2, something similar happens. 2) 2x - 11y = 6 (-3/2)*(2x) - 11(-3/2) = (-3/2)6 2**) -3x +(33/2)x = -9 Now compare 2**) with 1) 1) 3x - 9y = 6 2**) -3x +(33/2)x = -9 ----------------- Notice that the x variables in both equation are now easily eliminated by addition as with the first option. So, which is better? It is not clear. I think it may have to do with avoiding fractions.

Directrix (directrix):

Add all terms here and see how easily, or not, it is to solve for y. 1* with Equation 2 1*) -2x + 6y = -4 2) 2x - 11y = 6 -------------------- Then do the same with these two: 1) 3x - 9y = 6 2**) -3x + (33/2)x = -9 ------------------- @maaryliin Post what you decide, okay?